bill is 4 times as old as john.five year ago the sum of their ages was 50 how old are they now
Found 3 solutions by stanbon, Alan3354, TutorDelphia:
Answer by stanbon(75887)

bill is 4 times as old as john.five year ago the sum of their ages was 50 how old are they now
---
Equations:
b = 4*j
(b-5)+(j-5) = 50
-----
Substitute for "b" and solve for "j":
4j-5+j-5 = 50
5j = 60
j = 12 yrs. (john's age now)
----
Solve for "b":
b = 4*j
b = 4*12 = 48 yrs (Bill's age now)
=======================================

Answer by Alan3354(69443)

find the point (x,y) on the line y+-4x-5 that is equidistant from the points (8,-10) and (-7,0)
-------------
y+-4x-5 No equal sign

Answer by TutorDelphia(193)

bill is 4 times as old as john
b=4j
Five years ago bill was b-5 and john was j-5 (if bill is 48 now, he was 43, 5 years ago)
(b-5)+(j-5)=50
combine like terms
b+j-10=50
add 10 to both sides
b+j=60
sub in 4j for b
4j+j=60
combine like terms
5j=60
divide by 5
j=12
plug in 12 for j
b=48